Calculate the pH of a 0.59M solution of sodium acetate (NaCH3COO)


1
Expert's answer
2021-07-26T06:32:37-0400

Sodium acetate is a salt of a strong base and weak acid.

Ka of NaCH3COO "= 1.08 \times 10^{-5}"

"K_a \times K_b = 10^{-14} \\

K_b = \frac{10^{-14}}{1.8 \times 10^{-5}} = 5.56 \times 10^{-10}"


"K_b = 5.56 \times 10^{-10}= \frac{[CH_3COOH][NaOH]}{[CH_3COONa]} \\

= \frac{x \times x }{0.59-x} = \frac{x^2}{0.59-x}"

Since Kb is small, x <<<0.59

So, (0.59-x)≈0.59

"5.59 \times 10^{-10}= \frac{x^2}{0.59} \\

x = \sqrt{0.59 \times 5.56 \times 10^{-10}} = [OH^-] \\

[OH^-]=1.82 \times 10^{-5 } \\

pOH = -log[OH^-] = -log(1.82 \times 10^{-5})=4.74 \\

pH = 14-pOH \\

= 14-4.74 \\

= 9.26"

Answer: 9.26